OpenStudy (anonymous):

i'm having trouble converting radians to degrees: 7π/11 ty in adv

OpenStudy (anonymous):

Remember 2pi radians = 360 degrees... So I like writing these like this... \[(7\pi/11)/(2\pi) = x/360\] Which looks better when you write it out on paper. But then cross multiply and divide... \[(7\pi/11)*360= x*2\pi\] then just divide both by 2pi to get x.

OpenStudy (anonymous):

pi radians = 180 degrees so 7pi/11 = 7pi/11 / (180/ pi) degrees

OpenStudy (anonymous):

thanks to you both ( Tullee , jimmyrep ) for such fast replies these help a lot